OpenStudy (anonymous):

May I know what is e and ln,how to apply it to this 5^x=9

OpenStudy (e.mccormick):

e is a constant... it is like \(\pi\) in some respects. They represent numbers that never end. ln is shorthand for \(\log_e\) or log to the base of e. Because it is log base e, \(\ln e=1\) and \(\ln e^x\implies x \ln e=x\) and

OpenStudy (e.mccormick):

For \(5^x=9\) I would not use ln or e, but I would use logs. This talks about it: http://www.regentsprep.org/regents/math/algtrig/ATE8/exponentialEquations.htm The second set of examples, right at the top.

OpenStudy (anonymous):

So,when only should I use e and ln?

OpenStudy (e.mccormick):

Or not necessary use ln to be more clear. You can use log as well. For change of base stuff, like what your first one is, either ln or log will work. I would use whichever one your calculator does easiest. You need to use e if you are trying to solve ln of something: \(\ln x = 5\) \(e^{\ln x} = e^5\) \(x = e^5\)

OpenStudy (anonymous):

Ah,understood.Thanks :)

OpenStudy (e.mccormick):

np. And like I said, take a look at the examples on that page. They show it several times with solutions and steps. They short hand it, but it should help you to see them use it.

OpenStudy (e.mccormick):

Well, time to get some sleep... have fun!

OpenStudy (anonymous):

Alright,will do.